# Engagement Maximization

### v5

@@ 3.5 Illustrative Examples
is identical to the student-optimal solution (and independent to $\lambda$). This is an illustration of the conclusion of Proposition 5. The unknown parameter $\lambda$ can be pinned down by setting the student’s IC condition binding. The analysis above suggests that when the teacher only cares about the test-irrelevant knowledge, she will give the student exactly his preferred test-relevant information and extra knowledge such that the student gets barely enough welfare to be willing to participate.
@@ 4 Characterization of dynamics
In Section 3, we showed that to solve for the optimal policy, it is sufficient to study a simple static rational inattention problem (R) and implement the dynamic policy by diluting the static solution. In this section, we seek to find necessary conditions that characterize the dynamics of the optimal policy. For the purpose of characterizing optimal polices, for this section, we focus on the non-degenerate case where the optimal policy is to acquire some information ($\bar{J}(\bar{q}_{0})>0$ in the context of Proposition 1). We show that two desiderata—interim incentive compatibility and credibility—jointly provide a tight characterization of the belief dynamics.
@@ 4.1 Implications of interim incentive compatibility
$\overline{V}_{R}(q,\pi)=\max_{\pi^{\prime}\in\mathcal{P}(\text{Supp}(\pi)):E^{\pi^{\prime}}[q^{\prime}]=q}\mathbb{E}^{\pi^{\prime}}\left[\widehat{\mu}(q^{\prime})-\hat{u}(q)-\frac{\bar{\kappa}}{\chi}(H(q^{\prime})-H(q))\right].$
@@
We interpret $\overline{V}_{R}(q,\pi)$ as describing the maximum possible value of information acquisition, in the sense that $\overline{V}_{R}(q,\pi)$ is the difference between the best possible utility the agent could achieve with and without information acquisition, under the restriction that the set of stopping beliefs lie in the support of $\pi$ [^19] The value of information acquisition is related to, but not the same as, the notion of uncertainty defined in Frankel and Kamenica (2019). We define the set of “suspensive” beliefs given $\pi$ as the set for which the value of information acquisition is weakly positive.
@@
Suppose the optimal policy in (R), $\pi^{*}$, is unique. In this case, the stochastic process for beliefs $q_{t}$ cannot leave the set of suspensive beliefs before $t=\tau$. If $q_{t}$ left the convex hull of the support of $\pi^{*}$, then $\pi^{*}$ would not describe the law of $q_{\tau}$. If $q_{t}$ was decisive or if $\overline{V}_{R}(q_{t},\pi^{*})<0$, the agent would choose to stop. In the case of multiple optimal policies, this argument applies for some optimal $\pi^{*}$, which leads to the proposition below. Note that this proposition and the one that follows will include a technical caveat about finite support (that we believe is inessential), to simplify the proofs that follow. [^20] The assumption of finite support avoids the need to carefully define conditional expectations.
@@
Given $\bar{q}_{0}\in\mathcal{P}(X)$, let $\Pi^{*}$ be the set of solutions to (R). Suppose $(P,I^{H},\tau^{*})$ solves the principal’s problem under full commitment (Definition 1) and $\text{Supp}(q_{t}|t<\tau)$ has finite support for all $t$, then, $\forall t\in[0,\tau)$, $\forall q\in\text{Supp}(q_{t}|t<\tau)$, $q$ is suspensive given some $\pi$:
$\sup_{\pi\in\Pi^{*}}\overline{V}_{R}(q,\pi)\geq 0.$
@@
By Theorem 1, there is some $\pi^{*}\in\Pi^{*}$ that is the law of $q_{\tau^{*}}$ under the optimal policies, and the second claim follows. By the martingale property of the belief process, $q_{t}=E^{P^{*}}[q_{\tau}|\mathcal{F}_{t},t<\tau^{*}]$ for all $t\in[0,\tau^{*})$, and thus $q_{t}$ must lie in the convex hull of the support of $\pi^{*}$. The fact that $t<\tau^{*}$ implies $J(t)>0$ (i.e. $q_{t}\not\in\text{Supp}(\pi^{*})$), as otherwise, the principal cannot induce the agent to continue. Let $\pi^{\prime}$ be the distribution of $q_{\tau^{*}}$ conditional on $q_{t},t<\tau^{*}$. The agent’s incentive compatibility implies $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi^{\prime}}[H(q)-H(q_{t})]\leq\mathbb{E}^{\pi^{\prime}}[\widehat{\mu}(q)-\widehat{\mu}(q_{t})]$, from which it follows immediately that $\overline{V}_{R}(q_{t},\pi^{*})\geq 0$. ∎
@@
Implication 3: The Necessity of Jumps in Beliefs.
We have shown in Proposition 6 that, at all times under an optimal policy, beliefs are either suspensive or decisive. This result immediately implies that the optimal policy cannot involve continuous sample paths for beliefs (we assume the initial prior lies in the continuation region). If sample paths were continuous, the belief process $q_{t}$ would have to exit the set of suspensive beliefs, which would immediately induce the agent to stop. Consequently, beliefs must jump discontinuously from the set of suspensive beliefs (which lie in the continuation region of the benchmark model) to the set of decisive beliefs, which lie in the strict stopping region of the benchmark model (by Proposition 3).
@@ 4.2 Implications of credibility
So far, our analysis assumes full intertemporal commitment of the principal. However, in practice, commitment power might be limited, leading to a credibility concern. In this subsection, we tackle the issue of limited commitment by characterizing the necessary and sufficient conditions for a policy to be credible, i.e., immune to deviations in a sub-game perfection sense.
@@
$\forall q_{0}\in\mathcal{P}(X)$, $(P,I^{H},\tau)\in\mathcal{A}(q_{0})$ constitutes a strongly sub-game perfect equilibrium of the game if the agent’s interim incentive compatibility constraints are satisfied and
@@
    the principal would not deviate to the commitment solution: $\forall t$ and $q\in\text{Supp}(q_{t}|t\leq\tau)$, $\mathbb{E}^{P}[G(q_{\tau})-G(q_{t})|\mathcal{F}_{t},q_{t}=q]\geq J(q)$, and
@@
    the agent would not deviate by continuing under the agent-optimal benchmark: $\forall q\in\text{Supp}(q_{\tau})$, $V^{B}(q)\leq\hat{u}(q)$.
@@
In the analysis that follows, we will call a full-commitment optimal policy $(P,I^{H},\tau)$ credible if $(P,I^{H},\tau)$ is also a strongly sub-game perfect equilibrium in the game without commitment. The following theorem shows that the optimal dilution policy we construct in section 3 is indeed credible and hence constitutes a (strongly) sub-game perfect equilibrium. To avoid technicalities, what we will not discuss is whether there are other sub-game perfect equilibria that involve lower payoffs for the principal.
@@
$\forall\bar{q}_{0}\in\mathcal{P}(X)$, $(J(\bar{q}_{0},0))\in\mathcal{V}(\bar{q}_{0})$. Suppose $(P,I^{H},\tau)$ solves the principal’s problem under full commitment (Definition 1) and $\text{Supp}(q_{t}|t<\tau)$ has finite support for all $t$. $(P,I^{H},\tau)$ is credible if (and satisfies condition (i) of a strongly sub-game perfect equilibrium only if) the agent’s interim surplus is 00:
$\forall t\text{ and }q\in\text{Supp}(q_{t}|t<\tau),\quad\mathbb{E}^{P}[\hat{u}(q_{\tau})-\hat{u}(q)-\bar{\kappa}(\tau-t)|t<\tau]=0.$ (4)
@@
See Appendix section B.1. ∎
Note since $(J(\bar{q}_{0}),0)\in\mathcal{V}(\bar{q}_{0})$, $J(\bar{q}_{0})$ is itself a strongly sub-game perfect equilibrium payoff. Condition (i) in 3 must be satisfied for any weaker equilibrium notion that “selects SPNE in all sub-games in favor of the principal.” Therefore, (4) is a sufficient and necessary condition that guards the principal from deviations followed by a favorable selection of continuing equilibrium.
@@
$\underline{V}_{R}(q,\pi)=\min_{\pi^{\prime}\in\mathcal{P}(\text{Supp}(\pi)):E^{\pi^{\prime}}[q^{\prime}]=q}\mathbb{E}^{\pi^{\prime}}\left[\widehat{\mu}(q^{\prime})-\hat{u}(q)-\frac{\bar{\kappa}}{\chi}(H(q^{\prime})-H(q))\right].$
@@
Given $\pi\in\mathcal{P}(\mathcal{P}(X))$, the belief $q\in\mathcal{P}(X)$ is strictly suspensive if $q\ \text{ $\in$ Conv}(\text{Supp}(\pi))\setminus\mathrm{Supp(\pi)}$ and satisfies $\underline{V}_{R}(q,\pi)>0$.
@@
When $\pi^{\prime}=\pi$ is the only measure on the support of $\pi$ that satisfies Bayes-consistency, there is no distinction between $\bar{V}_{R}$ and $\underline{V}_{R}$. [^21] This will be true generically when the number of actions is weakly lower than the number of states. In this case, a belief is suspensive (resp. strictly suspensive) when this function is weakly (resp. strictly) positive.
@@
Given $\bar{q}_{0}\in\mathcal{P}(X)$, let $\Pi^{*}$ be the set of solutions to (R). Suppose $(P,I^{H},\tau)$ solves the principal’s problem under full commitment (Definition 1) and $\text{Supp}(q_{t}|t<\tau)$ has finite support for all $t$. $(P,I^{H},\tau)$ is credible only if $\forall t\in[0,\tau)$, $\forall q\in\text{Supp}(q_{t}|t<\tau)$, $q$ is not strictly suspensive given some $\pi$:
$\inf_{\pi\in\Pi^{*}}\underline{V}_{R}(q,\pi)\leq 0.$
@@
By Theorem 1, there is some $\pi^{*}\in\Pi^{*}$ that is the law of $q_{\tau^{*}}$ under the optimal policies. By the martingale property of the belief process, $q_{t}=E^{P^{*}}[q_{\tau}|\mathcal{F}_{t},t<\tau^{*}]$ for all $t\in[0,\tau^{*})$, and thus $q_{t}$ must lie in the convex hull of the support of $\pi^{*}$. Let $\pi^{\prime}$ be the distribution of $q_{\tau^{*}}$ conditional on $(q_{t},t<\tau^{*})$. Theorem 2 implies $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi^{\prime}}[H(q)-H(q_{t})]\geq\mathbb{E}^{\pi^{\prime}}[\widehat{\mu}(q)-\widehat{\mu}(q_{t})]$, from which it follows immediately that $\underline{V}_{R}(q_{t},\pi^{*})\leq 0$. ∎
Combining Propositions 6 and 7, in any solution that satisfies interim incentive compatibility and credibility, interim beliefs are suspensive but not strictly suspensive. They must be suspensive, as otherwise the agent would choose to stop, but cannot be strictly suspensive, as otherwise the principal would be tempted to deviate and extract more of the surplus.
Implication 4: Credible Beliefs Cannot Diffuse.
When beliefs are suspensive but not strictly suspensive, the agent’s value function is equal to $\hat{u}(q_{t})$ at all times. This implies that a pure diffusion process is infeasible almost everywhere in the continuation region. Because the agent recognizes that the principal will leave her with no surplus, only information that causes her to change her beliefs about the currently optimal action is valuable. If the belief $q_{t}$ is such that one action is strictly optimal given those beliefs (which will be true generically), the principal must offer at least the possibility of jumping to a different region in which another action is optimal; otherwise, the agent will perceive no benefit from the information provided. In the internet platform context, we interpret this result as suggesting that the principal provides the agent with news articles containing extreme or sensational claims, which should cause the agent to either move her beliefs a lot or not at all, as opposed to providing more nuanced or qualified information. In the teacher-student context, to engage a test-motivated student, the teacher must provide signals that have some possibility of changing the student’s belief about the correct answer on the test.
@@ 4.3 Illustrative Examples
We illustrate Propositions 6 and 7 by continuing example 1. First, focus on the $\bar{q}_{0}=\frac{1}{2}$ case. As is illustrated in Figure 1, the unique optimal $\pi^{*}$ involves two posterior beliefs $\left\{q^{1},q^{2}\right\}$, where $q^{1}<\underline{q}$ and $q^{2}>\bar{q}$. Per Proposition 2, the dilution of $\pi^{*}$ implements an optimal policy. In this case, the dilution of $\pi^{*}$ is the unique optimal policy. It is apparent from the symmetry of the problem that the value of information acquisition, $\overline{V}_{R}(q,\pi^{*})$, is maximized at $q=\frac{1}{2}$. As a result, the set of suspensive beliefs is a singleton, $\{\bar{q}_{0}\}$. By Proposition 6, beliefs will remain at $\bar{q}_{0}$ until they jump to either $q^{1}$ or $q^{2}$.
When $\bar{q}_{0}<\frac{1}{2}$, the optimal policy is not unique. In Figure 7, we plot the simplest dilution policy, where belief stays at $q_{0}$ until it jumps to the two optimal posteriors $q^{1}$ and $q^{2}.$ However, there can be other optimal policies. In this case, the set of suspensive beliefs is $[q_{0},1-q_{0}]$. Then, a full-commitment policy is optimal as long as beliefs remain in this interval $[q_{0},1-q_{0}]$ until eventually jumping to either $q^{1}$ or $q^{2}$. Specifically, Figure 7 plots one of such policies: when belief is at $q_{0}$ or $1-q_{0}$, it either jumps to the closer boundary, or drifts into the interior of $(q_{0},1-q_{0})$. When belief is interior, it follows a diffusion.
@@
We next illustrate the implications of Propositions 6 and 7 by continuing example 2, which discusses our two leading cases.
@@
Figure 9: Credible beliefs when $H=G$

### v6

@@ 3.3 Illustrative Examples
is identical to the student-optimal solution (solved in Example 1). This is an illustration of the conclusion of Proposition 5. The unknown parameter $\lambda$ can be pinned down by setting the student’s IC condition binding. The analysis above suggests that when the teacher only cares about the test-irrelevant knowledge, she will give the student exactly his preferred test-relevant information and extra knowledge such that the student gets barely enough welfare to be willing to participate.
@@ 4 Characterization of dynamics
In Section˜3, we showed that to solve for the optimal policy, it is sufficient to study a simple static rational inattention problem (R) and implement the dynamic policy by diluting the static solution. In this section, we seek to find necessary conditions that characterize the dynamics of the optimal policy. For the purpose of characterizing optimal polices, for this section, we focus on the non-degenerate case where the optimal policy is to acquire some information ($\bar{J}(\bar{q}_{0})>0$ in the context of Proposition 1). We show that two desiderata—interim incentive compatibility and credibility—jointly provide a tight characterization of the belief dynamics.
@@ 4.1 Implications of interim incentive compatibility
$\overline{V}_{R}(q,\pi)=\max_{\pi^{\prime}\in\mathcal{P}(\text{Supp}(\pi)):E^{\pi^{\prime}}[q^{\prime}]=q}\mathbb{E}^{\pi^{\prime}}\left[\widehat{u}(q^{\prime})-\widehat{u}(q)-\frac{\kappa}{\chi}(H(q^{\prime})-H(q))\right].$
@@
We interpret $\overline{V}_{R}(q,\pi)$ as describing the maximum possible value of information acquisition, in the sense that $\overline{V}_{R}(q,\pi)$ is the difference between the best possible utility the agent could achieve with and without information acquisition, under the restriction that the set of stopping beliefs lie in the support of $\pi$. [^21] The value of information acquisition is related to, but not the same as, the notion of uncertainty defined in frankel2019quantifying. We define the set of “suspensive” beliefs given $\pi$ as the set for which the value of information acquisition is weakly positive.
@@
Suppose the optimal policy in (R), $\pi^{*}$, is unique. In this case, the stochastic process for beliefs $q_{t}$ cannot leave the set of suspensive beliefs before $t=\tau$. If $q_{t}$ left the convex hull of the support of $\pi^{*}$, then $\pi^{*}$ would not describe the law of $q_{\tau}$. If $q_{t}$ was decisive or if $\overline{V}_{R}(q_{t},\pi^{*})<0$, the agent would choose to stop. In the case of multiple optimal policies, this argument applies for some optimal $\pi^{*}$, which leads to the proposition below.
@@
Given $\bar{q}_{0}\in\mathcal{P}(X)$, let $\Pi^{*}$ be the set of solutions to (R). Suppose $(P,\tau)$ solves (P) and $\sup_{\pi\in\Pi^{*}}\overline{V}_{R}(q,\pi)$ is a measurable function of $q$. Then, $\forall t\in[0,\tau)$, $q_{t}$ is almost surely suspensive:
$\mathrm{Prob}^{P}\left(\sup_{\pi\in\Pi^{*}}\overline{V}_{R}(q_{t},\pi)\geq 0\Big|t<\tau\right)=1.$
@@
See Appendix B.1. ∎
@@
Implication: The Necessity of Jumps in Beliefs.
A direct implication of Proposition 6 is that the optimal policy cannot involve continuous sample paths for beliefs (we assume the initial prior lies in the continuation region). If sample paths were continuous, the belief process $q_{t}$ would have to exit the set of suspensive beliefs, which would immediately induce the agent to stop. Consequently, beliefs must jump discontinuously from the set of suspensive beliefs (which lie in the continuation region of the benchmark model) to the set of decisive beliefs, which lie in the strict stopping region of the benchmark model (by Proposition 3).
To complement our discussion of the necessity of belief jumps, we analyze an extension of our model in which the belief process is required to be continuous. Formally, let $D^{C}\subset\mathbb{R}_{+}^{\mathcal{P}(X)}$ denote the set of continuous $\mathcal{P}(X)$-valued functions, i.e., all continuous paths. Let $\mathcal{A}^{C}(\bar{q}_{0})$ denote the subset of $\mathcal{A}(\bar{q}_{0})$ whose probability measures $P$ are supported within $D^{C}$. Consider the optimization problem:
$J^{C}(\bar{q}_{0})$ $=\sup_{(P,\tau)\in\mathcal{A}^{C}(\bar{q}_{0})}\mathbb{E}^{P}[G(q_{\tau})-G(\bar{q}_{0})|\mathcal{F}_{0}]$ (3)
s.t. $\tau\in\arg\max_{\tau^{\prime}\in\mathcal{T}}\mathbb{E}^{P}[\widehat{u}(q_{\tau^{\prime}})-\kappa\tau^{\prime}|\mathcal{F}_{0}].$
Equation (3) is the same as Definition 1 except that the set of admissible strategies is restricted to have no jumps.
We begin our analysis of this restricted problem by observing that the agent-optimal benchmark is unchanged. This follows from results in hebert2019rational, who show that the agent-optimal policy can be implemented by a pure diffusion process. Let $E^{A}=\left\{q\in\mathcal{P}(X)|V^{B}(q)>\widehat{u}(q)\right\}$ be the continuation region of the agent-optimal benchmark and $\bar{E}^{A}$ be the closure of $E^{A}$.
As we argued previously, the beliefs $q_{t}$ must lie in the continuation region of the agent-optimal benchmark if $t<\tau$. Given that the $\mathcal{A}^{C}(\bar{q}_{0})$ admits only continuous processes, it follows that the stopping beliefs must lie in $\bar{E}^{A}$.
Lemma 3.
If $(P,\tau)$ is admissible in Equation (3), then $\mathrm{Supp}(q_{\tau})\subset\bar{E}^{A}$.
Proof.
See the appendix, section B.2. ∎
If in addition there are only two states $(|X|=2$), the locally invariant posteriors property implies that the stopping beliefs will be identical to those the agent would choose in the agent-optimal problem; the only alternative the principal could choose would involve less information acquisition by the agent. [^22] When there are more than two states ($|X|>2$), the principal’s and agent’s optimal policies need not coincide. Consider as an example the case of two actions ($|A|=2$). The agent-optimal policy will in this case involve a diffusion on a line segment within the probability simplex (see hebert2019rational). The principal cannot induce the agent to follow this line segment beyond its endpoints, but can send the agent signals that cause the agent’s beliefs to move orthogonal to this line segment.
Proposition 7.
If $|X|=2$, then $J^{0}(\bar{q}_{0})$ is achieved by an agent-optimal policy.
Proof.
See the appendix, section B.3. ∎
This proposition illustrates that the possibility of jumps in beliefs is required to guarantee the agent’s participation constraint will bind. Without jumps in the belief process, the agent’s ability to stop at any time allows the agent to extract rents that would otherwise accrue to the principal.
@@ 4.2 Implications of credibility
So far, our analysis assumes full intertemporal commitment on the part of the principal. However, in practice, commitment power might be limited, leading to a credibility concern. In this subsection, we tackle the issue of limited commitment by characterizing the necessary and sufficient conditions for a policy to be credible, i.e., immune to deviations in a sub-game perfection sense.
@@
$\forall q_{0}\in\mathcal{P}(X)$, $(P,\tau)\in\mathcal{A}(\bar{q}_{0})$ constitutes a strongly sub-game perfect equilibrium of the game if the agent’s interim incentive compatibility constraints are satisfied and, in addition,
@@
    the principal would not deviate to the commitment solution: $\forall t$, $\mathrm{Prob}^{P}\Big(\mathbb{E}^{P}[G(q_{\tau})-G(q_{t})-J(q_{t})|\mathcal{F}_{t}]\geq 0\Big)=1$, and
@@
    the agent, when stopping, would not prefer to continue under the agent-optimal benchmark: $\mathrm{Prob}^{P}\big(V^{B}(q_{\tau})\leq\widehat{u}(q_{\tau})\big)=1$.
@@
In the analysis that follows, we will call a full-commitment optimal policy $(P,\tau)$ credible if $(P,\tau)$ is also a strongly sub-game perfect equilibrium in the game without commitment. The following theorem shows that the optimal dilution policy we construct in section˜3 is indeed credible and hence constitutes a (strongly) sub-game perfect equilibrium. To avoid technicalities, what we will not discuss is whether there are other sub-game perfect equilibria that involve lower payoffs for the principal. Let $\mathcal{P}(X)^{+}:=\{q\in\mathcal{P}(X)|G(q)<\sum q_{x}G(e_{x})\}$, namely, the set of beliefs that information has some value for the principal.
@@
$\forall\bar{q}_{0}\in\mathcal{P}(X)$, $J(\bar{q}_{0})$ can be achieved via a credible strategy, i.e., $(J(\bar{q}_{0}),0)\in\mathcal{V}(\bar{q}_{0})$. Suppose $(P,\tau)$ solves (P) and $\mathrm{supp}(q_{\tau})\subset\mathcal{P}(X)^{+}$. $(P,\tau)$ is credible if and only if the agent’s interim surplus is almost surely 0:
$\forall t,\quad\mathrm{Prob}^{P}\left(\mathbb{E}^{P}[\widehat{u}(q_{\tau})-\widehat{u}(q)-\kappa(\tau-t)|\mathcal{F}_{t}]=0\right)=1.$ (4)
@@
See Appendix B.4. ∎
Note since $(J(\bar{q}_{0}),0)\in\mathcal{V}(\bar{q}_{0})$, $J(\bar{q}_{0})$ is itself a strongly sub-game perfect equilibrium payoff. Condition (i) in ˜3 must be satisfied for any weaker equilibrium notion that “selects SPNE in all sub-games in favor of the principal.” Therefore, (4) is a sufficient and necessary condition that guards the principal from deviations followed by a favorable selection of continuing equilibrium.
@@
$\underline{V}_{R}(q,\pi)=\min_{\pi^{\prime}\in\mathcal{P}(\text{Supp}(\pi)):E^{\pi^{\prime}}[q^{\prime}]=q}\mathbb{E}^{\pi^{\prime}}\left[\widehat{u}(q^{\prime})-\widehat{u}(q)-\frac{\kappa}{\chi}(H(q^{\prime})-H(q))\right].$
@@
[…]
@@
When $\pi^{\prime}=\pi$ is the only measure on the support of $\pi$ that satisfies Bayes-consistency, there is no distinction between $\overline{V}_{R}$ and $\underline{V}_{R}$. [^23] This will be true generically when the number of actions is weakly lower than the number of states. In this case, a belief is suspensive (resp. strictly suspensive) when this function is weakly (resp. strictly) positive.
@@
Given $\bar{q}_{0}\in\mathcal{P}(X)$, let $\Pi^{*}$ be the set of solutions to (R). Suppose $(P,\tau)$ solves (P), $\mathrm{supp}(q_{\tau})\subset\mathcal{P}(X)^{+}$, and $\inf_{\pi\in\Pi^{*}}\underline{V}_{R}(q,\pi)$ is a measurable function of $q$. $(P,\tau)$ is credible only if $\forall t\in[0,\tau)$, $q_{t}$ is almost never strictly suspensive:
$\mathrm{Prob}^{P}\left(\inf_{\pi\in\Pi^{*}}\underline{V}_{R}(q,\pi)\leq 0\Big|t<\tau\right)=1.$
@@
See Appendix B.5. ∎
Combining Propositions 6 and 8, in any solution that satisfies interim incentive compatibility and credibility, interim beliefs are suspensive but not strictly suspensive. They must be suspensive, as otherwise the agent would choose to stop, but cannot be strictly suspensive, as otherwise the principal would be tempted to deviate and extract more of the surplus.
Implication: Credible Beliefs Cannot Diffuse.
When beliefs are suspensive but not strictly suspensive, the agent’s value function is equal to $\widehat{u}(q_{t})$ at all times. This implies that a pure diffusion process is infeasible almost everywhere in the continuation region. Because the agent recognizes that the principal will leave her with no surplus, only information that causes her to change her beliefs about the currently optimal action is valuable. If the belief $q_{t}$ is such that one action is strictly optimal given those beliefs (which will be true generically), the principal must offer at least the possibility of jumping to a different region in which another action is optimal; otherwise, the agent will perceive no benefit from the information provided. In the internet platform context, we interpret this result as suggesting that the principal provides the agent with news articles containing extreme or sensational claims, which should cause the agent to either move her beliefs a lot or not at all, as opposed to providing more nuanced or qualified information. In the teacher-student context, to engage a test-motivated student, the teacher must provide signals that have some possibility of changing the student’s belief about the correct answer on the test.
@@ 4.3 Illustrative Examples
We illustrate Propositions 6 and 8 by continuing example 1. First, focus on the $\bar{q}_{0}=\frac{1}{2}$ case. As is illustrated in Figure 1, the unique optimal $\pi^{*}$ involves two posterior beliefs $\left\{q^{1},q^{2}\right\}$, where $q^{1}<\underline{q}$ and $q^{2}>\bar{q}$. Per Proposition 2, the dilution of $\pi^{*}$ implements an optimal policy. In this case, the dilution of $\pi^{*}$ is the unique optimal policy. It is apparent from the symmetry of the problem that the value of information acquisition, $\overline{V}_{R}(q,\pi^{*})$, is maximized at $q=\frac{1}{2}$. As a result, the set of suspensive beliefs is a singleton, $\{\bar{q}_{0}\}$. By Proposition 6, beliefs will remain at $\bar{q}_{0}$ until they jump to either $q^{1}$ or $q^{2}$.
When $\bar{q}_{0}<\frac{1}{2}$, the optimal policy is not unique. In Figure 7, we plot the simplest dilution policy, where belief stays at $q_{0}$ until it jumps to the two optimal posteriors $q^{1}$ and $q^{2}.$ However, there can be other optimal policies. In this case, the set of suspensive beliefs is $q_{0},1-q_{0}]$. Then, a full-commitment policy is optimal as long as beliefs remain in this interval $q_{0},1-q_{0}]$ until eventually jumping to either $q^{1}$ or $q^{2}$. Specifically, Figure 7 plots one of such policies: when belief is at $q_{0}$ or $1-q_{0}$, it either jumps to the closer boundary, or drifts into the interior of $(q_{0},1-q_{0})$. When belief is interior, it follows a diffusion.
@@
We next illustrate the implications of Propositions 6 and 8 by continuing example 2, which discusses our two leading cases.
@@
Figure 9: Credible interim beliefs
when $H=G$